OpenStudy (anonymous):

solve each equation by graphing: x+y=3 and 2x-y=2 HELP PLEASE!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

x+y=3 x=-y+3 y=-x+3

OpenStudy (anonymous):

how did u get tht??

OpenStudy (anonymous):

from app :) i hope it's help you

OpenStudy (anonymous):

i mean how did you get x+y=3 x=-y+3 y=-x+3 ???????

OpenStudy (anonymous):

if you want to find x you put y in the other side and change it from +y to -y same goes with y.